OpenStudy (anonymous):

If the sinA = 2/5 and the cosA = 3/5, then what is tan A?

OpenStudy (phi):

tan A = sin A / cos A can you use this to answer the question ?

OpenStudy (anonymous):

tanA is = sinA/cosA so tanA =(2/5)/(3/5) flip and multiply to get 2/3

OpenStudy (anonymous):

b/a=2/3